# Multi-Agent Retrieval-Augmented Framework for Evidence Based Counterspeech Against Health Misinformation

### v1

@@ A.2.1 Annotation guidelines and process
Label 1: "Health Misinformation", If the post contains health-related
misinformation.
@@
Misinformation includes claims that contradict established medical research,
public health guidelines, or authoritative sources such as the World Health
Organization (WHO) and the Centers for Disease Control and Prevention (CDC).
@@
Despite the structured guidelines, some posts posed challenges in classification, particularly those containing satirical statements, ambiguous claims, or partially correct yet misleading information. In the first stage, the Cohen’s Kappa agreement rates between annotators were $\kappa\geq 0.58$, $\kappa\geq 0.67$, and $\kappa\geq 0.54$, indicating moderate to substantial agreement. In the second stage, the agreement rates were $\kappa\geq 0.69$, $\kappa\geq 0.48$, and $\kappa\geq 0.56$, also reflecting a mix of moderate and substantial agreement. To resolve annotation disagreements, the annotators engaged in group discussions and majority voting, ensuring that the final classification was well-supported by evidence. In cases where uncertainties persisted, the posts were cross-checked against fact-checking websites such as Snopes, HealthFeedback, and FactCheck.org to verify the accuracy of the information. This multi-step validation process ensured that the final dataset maintained a high level of accuracy and reliability.
@@
A.2.2 Human validation
@@ 3. Improve structure for effective communication (acknowledgment  correction
A.5 Annotation guidelines for factual accuracy and informativeness assessment
@@
A.6 Human evaluation guidelines for counterspeech comparison

### v2

@@ A.2.1 Annotation guidelines and process
Label 1: "Health Misinformation", if the post contains health-related misinformation.
@@
Misinformation includes claims that contradict established medical research, public
health guidelines, or authoritative sources such as the World Health Organization
(WHO), the Centers for Disease Control and Prevention (CDC), and the National
Institutes of Health (NIH).
@@
Despite the structured guidelines, some posts posed challenges in classification, particularly those containing satirical statements, ambiguous claims, or partially correct yet misleading information. In the first stage, the Cohen’s Kappa agreement rates between annotators were $\kappa\geq 0.58$, $\kappa\geq 0.67$, and $\kappa\geq 0.54$, indicating moderate to substantial agreement. In the second stage, the agreement rates were $\kappa\geq 0.69$, $\kappa\geq 0.48$, and $\kappa\geq 0.56$, also reflecting a mix of moderate and substantial agreement. To resolve disagreements in the annotations, the annotators participated in group discussions to ensure the final classification was well-supported by evidence. In cases where uncertainties persisted, the posts were cross-checked against fact-checking websites such as Snopes, HealthFeedback, and FactCheck.org to verify the accuracy of the information. This multi-step validation process ensured that the final dataset maintained a high level of accuracy and reliability.
@@
A.2.2 RoBERTa-large Implementation Details
We fine-tuned the roberta-large [^15] https://huggingface.co/FacebookAI/roberta-large model to classify Reddit posts as health misinformation or not, using the post as input. The model was trained on a class-balanced dataset with a maximum sequence length of 128, a batch size of 16, and a learning rate of 2e-5 for 4 epochs. We used the Adam optimizer and evaluated performance on a 20% test set.
Table 6 summarizes the classifier’s performance across both classes. The model achieved a weighted F1-score of 0.76.
Class Precision Recall F1-Score
Not Health Misinformation 0.71 0.84 0.77
Health Misinformation 0.82 0.68 0.74
Weighted Average 0.77 0.76 0.76
Table 6: Performance of the RoBERTa-large classifier on the human-annotated Reddit test set.
A.2.3 Human validation
@@ 3. Improve structure for effective communication (acknowledgment  correction
A.5 Evaluation Across Topics
Our dataset covers a range of health misinformation topics, including COVID-19, HIV, and Influenza. To assess the robustness of our approach across domains, we aggregated evaluation results by topic. As shown in Table 7, our Multi-Agent framework achieves consistently strong average performance across all three topics.
Method Category Politeness Relevance Informativeness Factual Accuracy
LLM w DP COVID-19 0.45 (0.27) 0.74 (0.17) 0.75 (0.14) 0.78 (0.21)
HIV 0.57 (0.31) 0.78 (0.12) 0.76 (0.06) 0.75 (0.27)
Influenza 0.42 (0.20) 0.81 (0.10) 0.75 (0.08) 0.70 (0.24)
Average 0.48 (0.26) 0.78 (0.13) 0.75 (0.09) 0.74 (0.24)
LLM w PE COVID-19 0.78 (0.20) 0.69 (0.12) 0.76 (0.06) 0.81 (0.24)
HIV 0.83 (0.17) 0.71 (0.12) 0.75 (0.00) 0.79 (0.17)
Influenza 0.86 (0.13) 0.73 (0.08) 0.76 (0.06) 0.71 (0.25)
Average 0.82 (0.17) 0.71 (0.11) 0.76 (0.04) 0.77 (0.22)
Static RAG COVID-19 0.75 (0.18) 0.67 (0.12) 0.78 (0.08) 0.89 (0.17)
HIV 0.81 (0.13) 0.70 (0.17) 0.76 (0.06) 0.68 (0.35)
Influenza 0.75 (0.16) 0.73 (0.10) 0.76 (0.06) 0.75 (0.26)
Average 0.77 (0.16) 0.70 (0.13) 0.77 (0.07) 0.77 (0.26)
Dynamic RAG COVID-19 0.84 (0.23) 0.62 (0.20) 0.70 (0.10) 0.88 (0.19)
HIV 0.91 (0.12) 0.70 (0.14) 0.73 (0.08) 0.84 (0.27)
Influenza 0.82 (0.23) 0.69 (0.12) 0.74 (0.10) 0.66 (0.26)
Average 0.86 (0.19) 0.67 (0.15) 0.72 (0.09) 0.79 (0.24)
Multi-Agent (Ours) COVID-19 0.92 (0.05) 0.68 (0.17) 0.78 (0.08) 0.84 (0.19)
HIV 0.86 (0.14) 0.73 (0.09) 0.78 (0.08) 0.91 (0.17)
Influenza 0.93 (0.14) 0.71 (0.10) 0.74 (0.06) 0.79 (0.15)
Average 0.90 (0.11) 0.71 (0.12) 0.77 (0.07) 0.85 (0.17)
Table 7: Topic-wise evaluation results across all systems for COVID-19, HIV, and Influenza. Standard deviation is reported in parentheses.
A.6 Annotation guidelines for factual accuracy and informativeness assessment
The primary objective is to determine how well the model-generated evaluation scores align with human judgments. We randomly select 100 samples, and three annotators independently evaluate the generated responses based on informativeness and factual accuracy on a 1-5 scale. Initially, the broad rating scale results in variability in annotator agreement. To address this, further discussions among the annotators are held, leading to a final label that reflects a consensus-driven evaluation. Once the final human labels are established, we compared them with our model evaluation results.
@@ A.6 Annotation guidelines for factual accuracy and informativeness assessment
A.7 Human evaluation guidelines for counterspeech comparison
We conduct a comparative human evaluation of two different response strategies: Response A (Multi-Agent w/o Refinement Agent) and Response B (Multi-Agent w Refinement Agent). The goal is to determine human preference and assess how the Refinement Agent enhances response quality.
@@ A.7 Human evaluation guidelines for counterspeech comparison
A.8 Overall Effectiveness Evaluation
To assess the overall quality of responses in correcting misinformation, we conducted a human evaluation comparing three systems: our Multi-Agent framework (Response A), Static RAG (Response B), and Dynamic RAG (Response C).
We asked three annotators to independently review 100 sampled misinformation posts along with the counterspeech responses generated by each system. Annotators were instructed to select the response that was most effective overall. Effectiveness was defined as how well the response refutes the misinformation and whether it is persuasive, credible, and well-structured.
Annotation Guideline
Your task is to compare three responses (Response A, Response B, and Response C)
that address a specific misinformation post. You will assess which response is
more effective based on the criteria below.
Evaluation Criteria:
1. Clarity: Is the response easy to understand? Does it avoid technical jargon
or overly complex language?
2. Politeness: Is the tone respectful and considerate? Does it correct
misinformation without sounding rude or confrontational?
3. Factual Accuracy: Are the facts correct and supported by evidence?
Does the response reference or align with reliable sources?
4. Informativeness: Does the response provide useful explanations, context,
or next steps? Is it educational and helpful in promoting understanding?
5. Overall Effectiveness: Does the response effectively refute misinformation?
Is it persuasive, credible, and well-structured?
How to Annotate:
1. Carefully read the Misinformation Post to understand the claim.
2. Compare Response A, Response B and Response C) across all five dimensions.
3. Choose the response that is more effective overall in correcting misinformation.